OpenStudy (anonymous):

Find the limit, if it exists. lim h -> 0, ((x+h)^3-x^3))/h

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{(x + h)^3 - x^3}{h}=\] \[\lim_{h \rightarrow 0} \frac{x^3 + 3x^2h+3xh^2+h^3 - x^3}{h}=\] \[\lim_{h \rightarrow 0} \frac{3x^2h+3xh^2+h^3}{h}=\] This should be easy to finish :)

OpenStudy (anonymous):

its not

OpenStudy (anonymous):

finish it

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \left(3x^2+3xh+h^2\right)=3x^2\]

OpenStudy (anonymous):

is that the finished product?

OpenStudy (anonymous):

its a multiple choice question but that isnt one of the answers i see

OpenStudy (anonymous):

What answers are there?

OpenStudy (anonymous):

A) 0 B) Does not exist C) 3x2 D) 3x2 + 3xh + h2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

ohh i see sorry

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok